Shift summary day 11 Jan. 19100


Subject: Shift summary day 11 Jan. 19100
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:19:33 CST


Shift summary day 11 Jan. 19100
(my computer back in chicago has a strange y2k bug so I
though I would repeat its form in the date for you here)

On Shift: RW, NS

Hours of beam available at all intensities: 4.0
Hours of beam available at nominal intensity: 3.5
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 3.0

Accounting of lost time:
o Tevatron quench
o Klystron problems, took advantage to chang over to short regenertor
run configuration.

========================================================================
Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: Leo and Sasha to fix stream 5 hangup problem.
                       Nick, Leo, Ed, Valerie and Rick K. to change
                           over to short regnertor run.
========================================================================
Runs ended this shift:
15478--special run for BA calibration with electrons.
15479--special run for BA calibration with electrons.

=======================================================================
Overview of shift:

Run Spill Time Comment
15478 34 07:00 New shift crew arrives, special run with TRD
                        offset and one hole pluged to be used for
                        electron identification by BA.
              07:07 TeV quench, no beam, estimate 1 hour.
              08:20 Sasha and Leo make access to fix stream 5.
              08:24 Repair man came to fix heater in gas shed.
              08:31 MCR message they are turning on TeV expect
                        beam in 30 min.
              08:50 Access done.
              09:05 MCR message says "trouble with tev high
                        order power supply", someone explain this
                        to me when and if you figure out what they
                        mean by "high order"
        164 09:54 Low intensity beam returns. DA still having
                        problems.
        184 10:20 Ending run to do a restart on DA to see if
                        that solves our problems.
15479 10:28 Run started, and no ktev ana erros for 2
                        spills!!!!!!!
              10:30 The maintance crew working on gas shead
                        heater comes in and says that he cannot fix
                        the heater until he shuts all the gas flow
                        in the building off and shut the electricity
                        to that room. We should call him when there is
                        a day or after the run is over. Lets hope the
                        weather does not freeze.
              10:35 Acceletor delivering nominal intensity.
          9 10:38 Beam gone.
         12 Beam back.
         20 beam gone
         23 beam back
         34 11:15 Beam gone, beam quench at D4, estimate 1
                        hour.
              11:20 Sasha put crates 16 and 17 (regenertor craates)
                        back into readout and DA still seams to work.
              11:45 Gas checklist done.
         82 12:15 Beam at low intensity returns.
       95-99 DA problems, Leo took crate 16 and 17 out of
                        readout and DA functions.
              12:45 Beam very intermitant since return, plus also
                        we took down the christmas tree lights..
        114 13:00 A spill at 9e12, requested intensity!!
                        Audio error stating L3 accept fraction error.
        129 Beam gone.
        131 Beam back.
              13:31 Leo called MCR and complained that the
                        beam spill structure was horrible and they
                        will look into it and see if the QXR needs
                        reset.
              13:50 Beam gone, many problems frig, linac and klystron.
        176 14:22 Run ended, MCR called to change over to
                        access mode.
              14:45 Access by Ed, Valerie, Nick, Rick K, and Leo:
                        to removed TRD from beam, ET delay cable and HA
                        threshold.
              14:50 Bob T. and Rick C. make access to NM2 target hall
                        to remove shadow absorber, beam plug and reconnect
                        regenerator cable.
              15:05 Regenerator moving started.
              15:15 NM2 out of access, returning keys to
                        MCR and should be ready to start run soon.
              15:30 End of shift, work to get short regenertor run
                        still in progress, need a muon run first then
                        move to the short regenerator run configuration.
        

        
=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:19:35 CST